# Quantum refraction effects in pulsar radio emission

### v1

@@ 3.2 For Case II
where $\hat{n}_{x\left[0\right]}$, $\hat{n}_{z\left[0\right]}$ and $\hat{n}_{y\left[0\right]}$ are given by equations (10), (11) and (26), respectively, and $B_{\mathrm{o}}=\mu\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}/r_{\mathrm{o}}^{3}$, and $\mathcal{I}\left(\lambda\right)$ refers to equation (30). In Fig. 4 is plotted the change in the polarization vector $\left(\Delta\varepsilon_{x},\Delta\varepsilon_{y},\Delta\varepsilon_{z}\right)\equiv\left.\left(\varepsilon_{x}\left(\lambda\right),\varepsilon_{y}\left(\lambda\right),\varepsilon_{z}\left(\lambda\right)\right)\right|_{0}^{\lambda}$ against $0\leq\lambda\leq 10$, wherein $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$ and $\eta{}_{1}B_{\mathrm{o}}^{2}\sim 10^{4}$ (fairly exaggerated for the purpose of visualization). From this plot, one can see that the polarization vector changes along the $x$-axis and $z$-axis while keeping the linear relationship between the components along these axes.
@@
Figure 4: The change in the polarization vector, $\left(\Delta\varepsilon_{x},\Delta\varepsilon_{y},\Delta\varepsilon_{z}\right)\equiv\left.\left(\varepsilon_{x}\left(\lambda\right),\varepsilon_{y}\left(\lambda\right),\varepsilon_{z}\left(\lambda\right)\right)\right|_{0}^{\lambda}$ plotted against $0\leq\lambda\leq 10$; the red solid curve represents the total polarization vector (classical polarization vector $+$ quantum correction), while the blue and green dashed curves represent the classical polarization vector and the quantum correction, respectively. For the purpose of visualization, the quantum refraction effect is fairly exaggerated by $\eta_{1}B_{\mathrm{o}}^{2}\sim 10^{4}$, $10^{9}$ times as large as an actual value $\sim 10^{-5}$, as in Fig. 3.
@@
The orthogonality between the propagation and polarization vectors,$\mathbf{\ \hat{n}}$ and $\mathbf{\varepsilon}$ can be inspected using equations (24)-(26) and (47)-(49):
@@
$\mathbf{\hat{n}}\cdot\mathbf{\varepsilon}$ $\approx\frac{4\eta_{1}B_{\mathrm{o}}^{2}\sin^{2}\left(\theta_{\mathrm{o}}-\alpha\right)\left(\hat{n}_{x\left[0\right]}^{2}-\hat{n}_{z\left[0\right]}^{2}\right)\hat{n}_{y\left[0\right]}\lambda^{2}}{\hat{n}_{x\left[0\right]}\hat{n}_{z\left[0\right]}\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)\left[\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\lambda^{2}+4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\lambda+\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\right]^{4}}$
@@
$\hskip 10.0pt-\frac{8\eta_{1}B_{\mathrm{o}}^{2}\Omega r_{\mathrm{o}}\sin^{2}\left(\theta_{\mathrm{o}}-\alpha\right)}{c\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)}\left(\hat{n}_{x\left[0\right]}+\frac{\sin\alpha}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}\right)\mathcal{I}\left(\lambda\right)+\mathcal{O}\left(\left(\Omega r_{\mathrm{o}}/c\right)^{2},_{\left[2\right]}\right),$ (50)
@@
where $\hat{n}_{x\left[0\right]}$, $\hat{n}_{z\left[0\right]}$ and $\hat{n}_{y\left[0\right]}$ are given by equations (10), (11) and (26), respectively, and $\mathcal{I}\left(\lambda\right)$ refers to equation (30). This implies that the quantum refraction effect results in breaking the orthogonality at the leading order in $\eta_{1}B_{\mathrm{o}}^{2}$. Note here that unlike Case I the departure from the orthogonality changes over the propagation of the light ray under this effect. For example, with $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$ and a usual value of $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$, we find the maximum departure value to be approximately $1.21\times 10^{-10}\,\mathrm{rad}$.
@@
$\varphi_{\mathrm{F}}=\cos^{-1}\left(\mathbf{\varepsilon}_{\left[0\right]}\cdot\mathbf{\hat{\varepsilon}}\right)\approx\sqrt{\delta\mathbf{\varepsilon}_{\left[1\right]}^{2}-\left(\mathbf{\varepsilon}_{\left[0\right]}\cdot\delta\mathbf{\varepsilon}_{\left[1\right]}\right)^{2}}+\mathcal{O}\left(\left(\Omega r_{\mathrm{o}}/c\right)^{2},_{\left[2\right]}\right).$ (51)
For example, with $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$ and a usual value of $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$, we find $\varphi_{\mathrm{F}}\approx 1.65\times 10^{-10}\,\mathrm{rad}$.
Note that the values, $\mathbf{\hat{n}}\cdot\mathbf{\varepsilon}\approx 1.21\times 10^{-10}\,\mathrm{rad}$ and $\varphi_{\mathrm{F}}\approx 1.65\times 10^{-10}\,\mathrm{rad}$, obtained as above are extremely small compared to our perturbation parameter $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$; therefore, these quantum refraction effects can be considered practically negligible.
@@ 3.3 Quantum birefringence
From equations (43) and (50) above, one can note the following: the polarization vector partly has a longitudinal component (i.e., a component parallel to the propagation vector) for Case I, whereas it is substantially perpendicular to the propagation vector for Case II. This is because given the conditions for pulsar emission as above, we have $\mathbf{\hat{n}}\cdot\mathbf{\varepsilon}_{\left(\mathrm{Case~{}I}\right)}\approx 3.02\times 10^{-5}\,\mathrm{rad}$, which is small but comparable to the perturbation parameter $\eta_{2}B_{\mathrm{o}}^{2}\approx 4.29\times 10^{-5}$, and therefore not negligible, while $\mathbf{\hat{n}}\cdot\mathbf{\varepsilon}_{\left(\mathrm{Case~{}II}\right)}\approx 1.21\times 10^{-10}\,\mathrm{rad}$ is practically negligible compared to the perturbation parameter $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$. The two different polarization modes, together with the dual refractive index $n$ as given by (31), are entirely due to the quantum refraction effect. These optical properties can be considered to define ‘quantum birefringence’ as the phenomenology involved is analogous to classical birefringence.
@@ 4 Conclusions and discussion
We have investigated the quantum refraction effects on the propagation and polarization of a photon in the dipole magnetic field background in pulsar radio emission, based on the PM Lagrangian from the HES one-loop action. Our main results are given by (24)-(26) and (27)-(29) (for the effects on the propagation) and by (40)-(42) and (47)-(49) (for the effects on the polarization); via perturbation analysis, we have determined the leading-order corrections to both the propagation and polarization vectors due to quantum refraction, which result in the deflection of the propagation (as shown by Fig. 2) and the change of the two polarization modes (as shown by Figs. 3 and 4). Further, we have inspected how quantum refraction affects the orthogonality between the propagation and polarization vectors and the Faraday rotation angle, the results of which are given in the texts at the end of Sections 3.1 and 3.2 for the two polarization modes. From the dual refractive index and the associated polarization modes under the effect of quantum refraction, we have discussed quantum birefringence, with the optical phenomenology analogous to classical birefringence.
@@ Appendix A Kinematic properties of a twisted curve and the Frenet–Serret formulas
where $\mathbf{N}_{\left[0\right]}$ denotes the unperturbed (classical) part of $\mathbf{N}$, taken from equation (56) with all the terms of $\delta n_{\left[1\right]}$ removed, and $\delta n_{\left[1\right]}$, $\hat{n}_{x\left[0\right]}$ and $\hat{n}_{z\left[0\right]}$ are given by equations (58), (10) and (11), respectively. Here we see that the curvature $\kappa$ is solely due to the quantum refraction effect, being led by the parameter $\eta_{2}B_{\mathrm{o}}^{2}$. In Fig. 5a is plotted the curvature $\kappa$ (dimensionless, multiplied by $r_{\mathrm{o}}$) against $0\leq\lambda\leq 10$ (with the substitution $s=r_{\mathrm{o}}\lambda$ in equation (60)), and in Fig. 5b is plotted the corresponding curve $\left(X/r_{\mathrm{o}},0,Z/r_{\mathrm{o}}\right)$, which is projected onto to the $xz$-plane. Here we assume $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$ and $\eta_{2}B_{\mathrm{o}}^{2}\sim 10^{4}$ (fairly exaggerated for the purpose of visualization). Note the two points marked for $\kappa_{\min}$ and $\kappa_{\max}$ in each plot; in particular, the trajectory is curved the most downward and the most upward at these points.

### v2

@@ 4.2 For Case II
where $\hat{n}_{x\left[0\right]}$, $\hat{n}_{z\left[0\right]}$ and $\hat{n}_{y\left[0\right]}$ are given by equations (15), (16) and (31), respectively, and $B_{\mathrm{o}}=\mu\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}/r_{\mathrm{o}}^{3}$, and $\mathcal{I}\left(\lambda\right)$ refers to equation (35). In Fig. 4 is plotted the change in the polarization vector $\left(\Delta\varepsilon_{x},\Delta\varepsilon_{y},\Delta\varepsilon_{z}\right)\equiv\left.\left(\varepsilon_{x}\left(\lambda\right),\varepsilon_{y}\left(\lambda\right),\varepsilon_{z}\left(\lambda\right)\right)\right|_{0}^{\lambda}$ against $0\leq\lambda\leq 10$ on a logarithmic scale, wherein $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$ and $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$. From this plot, one can see virtually no difference between the total polarization vector and the classical polarization vector; practically, no quantum refraction effect exists. However, the polarization vector changes along the $x$-axis and $z$-axis due to the rotational effect, as can be seen from equation (50).
@@
Figure 4: The change in the polarization vector, $\left(\Delta\varepsilon_{x},\Delta\varepsilon_{y},\Delta\varepsilon_{z}\right)\equiv\left.\left(\varepsilon_{x}\left(\lambda\right),\varepsilon_{y}\left(\lambda\right),\varepsilon_{z}\left(\lambda\right)\right)\right|_{0}^{\lambda}$ plotted against $0\leq\lambda\leq 10$ on a logarithmic scale; the red solid curve and the blue dashed curve represent the total polarization vector (classical polarization vector $+$ quantum correction) and the classical polarization vector, respectively.
@@
The orthogonality between the propagation and polarization vectors,$\mathbf{\ \hat{n}}$ and $\boldsymbol{\varepsilon}$ can be inspected using equations (29)-(31) and (52)-(54):
@@
$\mathbf{\hat{n}}\cdot\boldsymbol{\varepsilon}$ $\approx\frac{4\eta_{1}B_{\mathrm{o}}^{2}\sin^{2}\left(\theta_{\mathrm{o}}-\alpha\right)\left(\hat{n}_{x\left[0\right]}^{2}-\hat{n}_{z\left[0\right]}^{2}\right)\hat{n}_{y\left[0\right]}\lambda^{2}}{\hat{n}_{x\left[0\right]}\hat{n}_{z\left[0\right]}\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)\left[\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\lambda^{2}+4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\lambda+\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\right]^{4}}$
@@
$\hskip 10.0pt-\frac{8\eta_{1}B_{\mathrm{o}}^{2}\Omega r_{\mathrm{o}}\sin\alpha\sin^{2}\left(\theta_{\mathrm{o}}-\alpha\right)}{c\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{3/2}}\mathcal{I}\left(\lambda\right)+\mathcal{O}\left(\left(\Omega r_{\mathrm{o}}/c\right)^{2},_{\left[2\right]}\right),$ (55)
@@
where $\hat{n}_{x\left[0\right]}$, $\hat{n}_{z\left[0\right]}$ and $\hat{n}_{y\left[0\right]}$ are given by equations (15), (16) and (31), respectively, and $\mathcal{I}\left(\lambda\right)$ refers to equation (35). This implies that the quantum refraction effect results in breaking the orthogonality at the leading order in $\eta_{1}B_{\mathrm{o}}^{2}$. Note here that unlike Case I the departure from the orthogonality changes over the propagation of the light ray under this effect. For example, with $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$ and a usual value of $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$, we find the maximum departure value to be $\sin^{-1}\left(\mathbf{\hat{n}}\cdot\boldsymbol{\varepsilon}\right)\approx 1.05\times 10^{-10}\,\mathrm{rad}$.
@@
$\varphi_{\mathrm{F}}=\cos^{-1}\left(\boldsymbol{\varepsilon}_{\left[0\right]}\cdot\boldsymbol{\hat{\varepsilon}}\right)\approx\sqrt{\delta\boldsymbol{\varepsilon}_{\left[1\right]}^{2}-\left(\boldsymbol{\varepsilon}_{\left[0\right]}\cdot\delta\boldsymbol{\varepsilon}_{\left[1\right]}\right)^{2}}+\mathcal{O}\left(\left(\Omega r_{\mathrm{o}}/c\right)^{2},_{\left[2\right]}\right).$ (56)
For example, with $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$ and a usual value of $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$, we find $\varphi_{\mathrm{F}}\approx 1.74\times 10^{-10}\,\mathrm{rad}$.
Note that the values, $\sin^{-1}\left(\mathbf{\hat{n}}\cdot\boldsymbol{\varepsilon}\right)\approx 1.05\times 10^{-10}\,\mathrm{rad}$ and $\varphi_{\mathrm{F}}\approx 1.74\times 10^{-10}\,\mathrm{rad}$, obtained as above are extremely small compared to our perturbation parameter $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$; therefore, these quantum refraction effects can be considered practically negligible.
@@ 4.3 Quantum birefringence
From equations (48) and (55) above, one can note the following: the polarization vector partly has a longitudinal component (i.e., a component parallel to the propagation vector) for Case I, whereas it is substantially perpendicular to the propagation vector for Case II. This is because given the conditions for pulsar emission as above, we have $\sin^{-1}\left(\mathbf{\hat{n}}\cdot\boldsymbol{\varepsilon}_{\left(\mathrm{Case~{}I}\right)}\right)\approx 3.02\times 10^{-5}\,\mathrm{rad}$, which is small but comparable to the perturbation parameter $\eta_{2}B_{\mathrm{o}}^{2}\approx 4.29\times 10^{-5}$, and therefore not negligible, while $\sin^{-1}\left(\mathbf{\hat{n}}\cdot\boldsymbol{\varepsilon}_{\left(\mathrm{Case~{}II}\right)}\right)\approx 1.05\times 10^{-10}\,\mathrm{rad}$ is practically negligible compared to the perturbation parameter $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$. The two different polarization modes, together with the dual refractive index $n$ as given by (36), are entirely due to the quantum refraction effect. These optical properties can be considered to define ‘quantum birefringence’ as the phenomenology involved is analogous to classical birefringence.
@@ 5 Conclusions and discussion
We have investigated the quantum refraction effects on the propagation and polarization of a photon in the dipole magnetic field background in pulsar emission, based on the PM Lagrangian from the HES one-loop action. Our main results are given by (29)-(31) and (32)-(34) (for the effects on the propagation) and by (45)-(47) and (52)-(54) (for the effects on the polarization); via perturbation analysis, we have determined the leading-order corrections to both the propagation and polarization vectors due to quantum refraction, which result in the deflection of the propagation (as shown by Fig. 2) and the change of the two polarization modes (as shown by Figs. 3 and 4). Further, we have inspected how quantum refraction affects the orthogonality between the propagation and polarization vectors and the Faraday rotation angle, the results of which are given in the texts at the end of Sections 4.1 and 4.2 for the two polarization modes. From the dual refractive index and the associated polarization modes under the effect of quantum refraction, we have discussed quantum birefringence, with the optical phenomenology analogous to classical birefringence.
@@
It should be noted that not only the quantum refraction due to strong fields but also the plasmas spreading around the magnetosphere might affect the propagation and polarization of radiation. In a broad spectrum of curvature radiation, the low-frequency part, e.g., radio emission, is dominated by the plasma effects, while the high-frequency part, e.g., X-ray emission, is dominated by the quantum refraction effects, as is well discussed in Wang & Lai (2007). To take the plasma effects into account properly, a good plasma distribution model is required, which is hardly available except for specific computational simulations (Petri (2016)). In contrast, our analysis for the quantum refraction effects is well-defined given a specific configuration of coherent magnetic fields (Kim & Trippe (2021)). Therefore, for observation of quantum refraction effects, X-ray emission is more favoured than other emissions in the lower frequency regime; for example, a space telescope (the Imaging X-ray Polarimetry Explorer) performed the X-ray polarimetry ($\sim$ a few $\mathrm{keV}$) to confirm the quantum refraction effects for a magnetar (Taverna et al. (2022)).
@@ Appendix A Kinematic properties of a twisted curve and the Frenet–Serret formulas
where $\mathbf{N}_{\left[0\right]}$ denotes the unperturbed (classical) part of $\mathbf{N}$, taken from equation (61) with all the terms of $\delta n_{\left[1\right]}$ removed, and $\delta n_{\left[1\right]}$, $\hat{n}_{x\left[0\right]}$ and $\hat{n}_{z\left[0\right]}$ are given by equations (63), (15) and (16), respectively. Here we see that the curvature $\kappa$ is solely due to the quantum refraction effect, being led by the parameter $\eta_{2}B_{\mathrm{o}}^{2}$. In Fig. 5a is plotted the curvature $\kappa$ (dimensionless, multiplied by $r_{\mathrm{o}}$) against $0\leq\lambda\leq 10$ (with the substitution $s=r_{\mathrm{o}}\lambda$ in equation (65)), and in Fig. 5b is plotted the corresponding curve $\left(X/r_{\mathrm{o}},0,Z/r_{\mathrm{o}}\right)$, which is projected onto to the $xz$-plane. Here we assume $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$ and $\eta_{2}B_{\mathrm{o}}^{2}\sim 10^{4}$ ($10^{9}$ times as large as an actual order $\sim 10^{-5}$; fairly exaggerated for intuitive visualization). Note the two points marked for $\kappa_{\min}$ and $\kappa_{\max}$ in each plot; in particular, the trajectory is curved the most downward and the most upward at the former and the latter points, respectively.